OpenStudy (anonymous):

Solve x2 + 8x – 48 = 0 by completing the square. Show your work for full credit.

OpenStudy (radar):

For completing the square the coefficient of the "squared" term should be 1 and it is.

OpenStudy (radar):

The term to be added to make it a perfect square is one half of the coefficent of the x term squared.

OpenStudy (radar):

What you add you must also subtract to retain the equations original value.

OpenStudy (anonymous):

\[(x^2+8x+16)-16-48\] <---- just added and substracted 4, which is \[(8/2)^2\] Now try to factor out the term in brackets what do you get?

OpenStudy (radar):

Lets see how saljudieh07 did this

OpenStudy (radar):

Looks good.

OpenStudy (anonymous):

sorry \[4^2\] that is what I added and substracted

OpenStudy (anonymous):

now, kaymarie12479 can you factor the expression in the brackets?

OpenStudy (radar):

Actually you want to the square root of the value in the brackets. The origiinal equation is factorable, but you want to "complete the square. \[(x + 4)^{2}=64\]

OpenStudy (radar):

\[x+4=\pm8\]

OpenStudy (radar):

Note the original equation is factorable (x-4)(x+12)=0 however, they want you to use completing the square method.

OpenStudy (radar):

Note. That factoring gives you x=4 and x=-12 as answer and x + 4 = +/- 8 gives same results x=4, and x=-12

OpenStudy (radar):

@saljudieh07 you left off the "=0" and disregarded your constants.